# Training-Free Text-Guided Color Editing with Multi-Modal Diffusion Transformer

### v2

@@ 4.1. Setup
Table 2. Quantitative image results compared with commercial models on PIE-Bench.
Benchmark. While prior editing methods (Hertz et al., 2023; Cao et al., 2023; Cai et al., 2025) typically lack standardized benchmark evaluation, we adopt prompts from the Change Color task in PIE-Bench (Ju et al., 2024), which includes 40 editing pairs, to better showcase the capabilities of our method. Although our approach is fully compatible with inversion methods, we adopt a noise-to-image setting to better isolate and evaluate editing performance, removing the influence of reconstruction and inversion. This also allows us to reuse the same set of benchmark prompts for video diffusion models, enabling a consistent and fair comparison across both image and video domains. To further scale up the evaluation, we introduce a new benchmark called ColorCtrl-Bench, consisting of 300 prompt pairs in the same style as PIE-Bench (see Sec. A.4 for details). For all baselines, we adopt a fixed sampler and identical random seeds to ensure that source images are consistent, enabling reliable comparison across methods.
Evaluation protocol. Unlike the original PIE-Bench, which evaluates structural similarity using structural distance (Tumanyan et al., 2022), we adopt the Structural Similarity Index (SSIM) (Wang et al., 2004) computed on Canny edge maps (Canny, 1986), following the approach of Zhao et al. (2023), for more accurate assessment. To evaluate the preservation of non-edited regions (a.k.a., BG Preservation), we compute PSNR and SSIM exclusively on those regions, which are annotated using Grounded SAM 2 (Ren et al., 2024) with dilation. Semantic alignment of the edits is assessed using CLIP similarity (Radford et al., 2021), applied to both the entire image and the edited regions.
4.2. Comparison with Training-Free Methods (Images)
4.3. Comparison with Commercial Models (Images)
@@
4.4. Comparison with Training-Free Methods (Videos)
   Method    PIE-Bench (Ju et al., 2024)    ColorCtrl-Bench
   Fix seed    0.6228    20.73    0.8912    27.97    26.63    0.6671    19.49    0.8489    28.01    26.66
   FireFlow (Deng et al., 2025)    0.7517    30.74    0.9605    25.42    24.18    0.7957    28.78    0.9323    25.26    24.12
   RF-Solver (Wang et al., 2025)    0.7677    35.41    0.9730    25.40    24.08    0.8102    32.64    0.9562    25.36    24.16
   SDEdit (Meng et al., 2022)    0.7880    35.46    0.9689    24.64    23.21    0.8281    34.30    0.9626    24.74    23.46
   DiTCtrl (Cai et al., 2025)    0.6912    27.70    0.9435    27.02    25.57    0.7270    26.14    0.9165    26.54    25.14
   UniEdit-Flow (Jiao et al., 2025)    0.7645    38.00    0.9772    25.35    24.13    0.8002    35.33    0.9613    25.39    24.14
   Ours    0.8651    38.98    0.9916    27.12    25.96    0.8885    38.68    0.9893    27.32    26.16
@@ 4.4. Comparison with Training-Free Methods (Videos)
Table 3. Quantitative video results compared with baselines on PIE-Bench and ColorCtrl-Bench.
Figure 5. Qualitative video results compared with training-free methods on PIE-Bench. Prompt is “green tea” $\to$ “red tea”. Each shows three frames.
@@
Thanks to the training-free nature of our method, it can be seamlessly extended to video editing tasks. Since FLUX.1 Kontext Max and GPT-4o do not support video editing, we compare only with other training-free methods. FlowEdit is excluded as it only applies to rectified flow models rather than diffusion-based models. Tab. 3 presents benchmark results on CogVideoX-2B (Yang et al., 2024). Similar to image editing, our method outperforms all baselines. Notably, the performance gap becomes even more pronounced due to the added temporal dimension. Visualization results in Fig. 5 further highlight the effectiveness of ColorCtrl. Consistent with its performance in image editing, our method also handles challenging cases in video, such as accurately reflecting the color change of the ice cream in the bowl’s reflection. Please refer to Appendix B for additional comparisons.
4.5. Attribute Re-Weighting Analysis
@@
Figure 6. Examples of attribute re-weighting. The top two rows are generated using FLUX.1-dev, while the bottom one are generated using SD3.
@@
Although DiTCtrl (Cai et al., 2025) includes a mechanism for attribute re-weighting, both the paper and results in Tab. 1 suggest limitations. Specifically, the method struggles to achieve the intended color changes while maintaining consistency with the source image. Moreover, their approach of scaling attention weights after the softmax operation violates the assumption that attention scores should sum to one, which leads to incorrect attention behavior. Since none of the baselines can smoothly adjust attribute strength while simultaneously satisfying the three constraints ((C1)-(C3)), we present results of our method alone in Fig. 6. These results show that our method not only supports re-weighting a single attribute within the same prompt, but also allows adjusting attribute strength across different prompts (second row). Moreover, our method can re-weight multiple attributes simultaneously (third row). Overall, these results demonstrate that our method enables smooth and controllable transitions in attribute strength, while preserving structural consistency across the image and maintaining color fidelity in non-edited regions, on both SD3 and FLUX.1-dev.
4.6. Ablation Study
@@
Figure 7. Examples of results generated with Step1X-Edit (left) and FLUX.1 Kontext dev (right). Green arrows: first edit using the editing model. Blue arrows: second edit using the editing model. Orange arrows: second edit using the editing model with ColorCtrl. Top left: a red diamond is added to the neck, then changed to blue. Bottom left: an orange cap is added, then changed to purple. Top right: blue butterflies are added, then changed to yellow. Bottom right: a flashlight with blue light is added, then changed to green.
@@
Table 4. Ablation study evaluating the effectiveness of each component on PIE-Bench.
4.7. Compatibility and Discussion
Real image editing.
Figure 8. Examples of real image editing. Left: results generated with SD3. Right: results generated with FLUX.1-dev.
To apply ColorCtrl to real images, we integrate UniEdit-Flow (Jiao et al., 2025) for inversion and replace the original editing module with our method. The results, shown in Fig. 8, demonstrate that our approach generalizes well to real-world inputs on both SD3 and FLUX.1-dev, preserving fine-grained consistency (e.g., subtle fabric wrinkles and shadows) while delivering strong editing performance. Notably, in the top row, even when editing black clothing, ColorCtrl accurately distinguishes material shading from cast shadows, resulting in illumination-consistent edits.
Generalization to instruction-based editing diffusion models. In addition to text-to-image and text-to-video models, ColorCtrl is also compatible with instruction-based editing diffusion models, such as Step1X-Edit (Liu et al., 2025a) and FLUX.1 Kontext dev (Black Forest Labs, 2025). Given a real input image and a target editing instruction, the model performs edits accordingly. However, performing a second round of editing using the original model alone often leads to structural inconsistencies, such as distortions or shifts in shadows and edges. By incorporating our method, the model can further refine color edits while preserving structural fidelity. As shown in Fig. 7, compared to using the base model alone, our approach achieves better outline consistency and improved preservation of subtle visual cues like shadows.
Discussion. Our method integrates structure preservation, regional color preservation, and word-level attribute intensity control into a unified, training-free pipeline. These components work together to enable precise and consistent text-driven color editing. Each part of the attention computation plays a distinct role: the vision-to-vision part of $\bm{M}$ preserves structure, the vision-to-text part is used for mask extraction, the text-to-vision part enables attribute re-weighting, the vision part of $\bm{V}$ supports color preservation, and the text-to-text region of $\bm{M}$ along with the text part of $\bm{V}$ provides crucial textual guidance. As demonstrated by the degradation observed in Fig. 3 (a) when the text-to-text region of $\bm{M}$ or the text part of $\bm{V}$ is altered, preserving the integrity of these parts is essential for maintaining robust and coherent generation that aligns with the target prompt.
5. Conclusion
@@ A.1. Inference Settings
The target object for mask extraction is determined using the “blended_word” keywords provided by PIE-Bench (Ju et al., 2024) and ColorCtrl. We use a mask threshold of $\epsilon=0.1$, which consistently yields strong performance across SD3, FLUX.1-dev, and CogVideoX-2B. Despite being relatively coarse, this threshold is sufficient due to the strong global adaptation ability of modern generative models, which enables them to propagate edits from sparse color cues to semantically aligned regions.
For Step1X-Edit (Liu et al., 2025a) and FLUX.1 Kontext dev (Black Forest Labs, 2025), we use their official code and setting with resolution 1024 $\times$ 1024.
@@
A.2. Sampling Details
@@
Table 5. Quantitative image results compared with training-free methods on ColorCtrl-Bench. Results for FireFlow on SD3 are omitted due to consistency worse than that obtained using fixed seeds.
A.3. Implementation of Compared Methods
Figure 9. Examples of FireFlow and RF-Solver with difference end timesteps on SD3. The selected end timestep refers to the setting used in the benchmark evaluation, while the higher end timestep denotes a larger value chosen for comparison purposes.
  • FireFlow (Deng et al., 2025): In SD3, we observe that it is difficult to select a suitable end timestep for FireFlow, as setting it too high often leads to artifacts and generation failures, as shown in Fig. 9. To mitigate quality degradation caused by excessive editing steps, we restrict editing to timesteps 0 through 3 across all blocks in SD3. In contrast, FLUX.1-dev does not exhibit this issue, and using the same range (timesteps 0 to 3) yields results comparable to those reported in the original paper. For video generation, the editing ends at timestep $9$. During editing, value tokens ($\bm{V}$) are copied from the source to the target.
A.4. ColorCtrl-Bench Construction
Here, we provide details on the construction of ColorCtrl-Bench. Following a similar approach to that in Ju et al. (2024), we use GPT to generate a dataset of tuples, each consisting of a source prompt, a target prompt, a subject token, and an instruction, consistent with the format used in PIE-Bench. These tuples represent an image before and after a color editing operation, along with an instruction describing the transformation and a subject token indicating the object to be modified. The exact template is shown in Fig. 17.
A.5. Evaluation Details
Appendix B More Results and Analysis
B.1. More Results of Image Editing
@@
Table 6. Quantitative image results compared with commercial models on ColorCtrl-Bench.
Figure 10. Examples of image editing results.
@@
B.2. More Results of Video Editing
@@
Fig. 14 and Fig. 11 present video editing results on ColorCtrl-Bench, while Fig. 16 shows additional results on PIE-Bench. Our method consistently outperforms all baselines, achieving superior color editing accuracy and better structural consistency across frames.
Figure 11. Additional qualitative video results compared with training-free models on ColorCtrl-Bench. Left: “gray sofa” $\to$ “brown sofa”. Right: “brown suitcase” $\to$ “black suitcase”. Each shows three frames.
B.3. Limitations

### v3

@@ 4.2 Comparison with Training-Free Methods (Images)
Figure 4: Qualitative image results compared with training-free methods and commercial models on PIE-Bench. The top three rows are generated using FLUX.1-dev, while the bottom two are generated using SD3. Best viewed with zoom-in.
4.3 Comparison with Commercial Models (Images)
@@
4.4 Comparison with Training-Free Methods (Videos)
@@
Table 3: Quantitative video results compared with baselines on PIE-Bench.
Method Canny BG Preservation $\text{Clip Similarity}$$\uparrow$
$\text{SSIM}$$\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ Whole Edited
Fix seed 0.6228 20.73 0.8912 27.97 26.63
FireFlow (Deng et al., 2025b) 0.7517 30.74 0.9605 25.42 24.18
RF-Solver (Wang et al., 2025) 0.7677 35.41 0.9730 25.40 24.08
SDEdit (Meng et al., 2022) 0.7880 35.46 0.9689 24.64 23.21
DiTCtrl (Cai et al., 2025) 0.6912 27.70 0.9435 27.02 25.57
UniEdit-Flow (Jiao et al., 2025) 0.7645 38.00 0.9772 25.35 24.13
Ours 0.8651 38.98 0.9916 27.12 25.96
@@
Thanks to the training-free nature of our method, it can be seamlessly extended to video editing tasks. Since FLUX.1 Kontext Max and GPT-4o do not support video editing, we compare only with other training-free methods. FlowEdit is excluded as it only applies to rectified flow models rather than diffusion-based models. Tab. 3 and Tab. 8 presents benchmark results on CogVideoX-2B (Yang et al., 2024). Similar to image editing, our method outperforms all baselines. Notably, the performance gap becomes even more pronounced due to the added temporal dimension. Visualization results in Fig. 6 further highlight the effectiveness of ColorCtrl. Consistent with its performance in image editing, our method also handles challenging cases in video, such as accurately reflecting the color change of the ice cream in the bowl’s reflection. Please refer to Appendix B for additional comparisons.
4.5 Attribute Re-Weighting Analysis
@@
Figure 5: Examples of attribute re-weighting. The top two rows are generated using FLUX.1-dev, while the bottom one are generated using SD3.
@@
Although DiTCtrl (Cai et al., 2025) includes a mechanism for attribute re-weighting, both the paper and results in Tab. 1 suggest limitations. Specifically, the method struggles to achieve the intended color changes while maintaining consistency with the source image. Moreover, their approach of scaling attention weights after the softmax operation violates the assumption that attention scores should sum to one, which leads to incorrect attention behavior. Since none of the baselines can smoothly adjust attribute strength while simultaneously satisfying the three constraints ((C1)-(C3)), we present results of our method alone in Fig. 5. These results show that our method not only supports re-weighting a single attribute within the same prompt, but also allows adjusting attribute strength across different prompts (second row). Moreover, our method can re-weight multiple attributes simultaneously (third row). Overall, these results demonstrate that ColorCtrl enables smooth and controllable transitions in attribute strength, while preserving structural consistency across the image and maintaining color fidelity in non-edited regions, on both SD3 and FLUX.1-dev.
4.6 Ablation Study
Figure 6: Qualitative video results compared with training-free methods on PIE-Bench. The edit prompt is “green tea” $\to$ “red tea”. Each shows three frames.
@@ 4.6 Ablation Study
4.7 Compatibility and Discussion
Figure 7: Examples of real image editing. Results generated with SD3 (left) and FLUX.1-dev (right).
Real image editing.
To apply ColorCtrl to real input images, we integrate the inversion method from UniEdit-Flow (Jiao et al., 2025) and replace the original editing module with our method. As shown in Fig. 7, our approach generalizes well to real-world inputs and matches its noise-to-image performance on both SD3 and FLUX.1-dev, preserving fine-grained consistency (e.g., subtle fabric wrinkles and shadows) while delivering strong editing performance. Notably, in the top row, even when editing black clothing, our method accurately distinguishes material shading from cast shadows, resulting in illumination-consistent edits.
Generalization to instruction-based editing diffusion models.
Figure 8: Examples of results generated with Step1X-Edit (left) and FLUX.1 Kontext dev (right). Green arrows: first edit using the editing model. Blue arrows: second edit directly using the editing model conditioned on the result of the first edit. Orange arrows: second edit using the editing model with ColorCtrl. Top left: a red diamond is added to the neck, then changed to blue. Bottom left: an orange cap is added, then changed to purple. Top right: blue butterflies are added, then changed to yellow. Bottom right: a flashlight with blue light is added, then changed to green.
In addition to text-to-image and text-to-video models, ColorCtrl is also compatible with instruction-based editing diffusion models, such as Step1X-Edit (Liu et al., 2025b) and FLUX.1 Kontext dev (Black Forest Labs, 2025). Given a real input image and a target editing instruction, the model performs edits accordingly. However, performing a second round of editing using the original model alone often leads to structural inconsistencies, such as distortions or shifts in shadows and edges. By incorporating our method, the model can further refine color edits while preserving structural fidelity in both edited and non-edited regions. As shown in Fig. 8, compared to using the base model alone, our approach achieves better outline consistency and improved preservation of subtle visual cues like shadows.
Additional results.
Appendix B presents benchmark evaluations on ColorCtrl-Bench, limitation analysis, user studies, and downstream applications.
Discussion.
Our method integrates structure preservation, regional color preservation, and word-level attribute intensity control into a unified, training-free pipeline. These components work together to enable precise and consistent text-driven color editing. Each part of the attention computation plays a distinct role: the vision-to-vision part of $\bm{M}$ preserves structure, the vision-to-text part is used for mask extraction, the text-to-vision part enables attribute re-weighting, the vision part of $\bm{V}$ supports color preservation, and the text-to-text region of $\bm{M}$ along with the text part of $\bm{V}$ provides crucial textual guidance. As demonstrated by the degradation observed in Fig. 3 (a) when the text-to-text region of $\bm{M}$ or the text part of $\bm{V}$ is altered, preserving the integrity of these parts is essential for maintaining robust and coherent generation that aligns with the target prompt.
Table 4: Ablation study evaluating the effectiveness of each component on PIE-Bench.
@@ Discussion.
5 Conclusion
@@
Ethics Statement
Advances in image editing technologies bring not only new capabilities but also ethical challenges. While our method improves the precision of color editing via text-based control, it may also be misapplied to produce deceptive or harmful visual content. To mitigate such risks, we advocate for responsible usage practices that prioritize transparency, user awareness, and informed consent in real-world applications. Moreover, since our method relies on pretrained models, it inherits any biases embedded in these models. Such biases may surface in the edited outputs in subtle or unintended ways. We consider this an open area for further investigation and support ongoing efforts to identify and reduce algorithmic bias. All human studies were conducted with voluntary participants who were informed of the goals of the study and provided explicit consent.
Reproducibility Statement
To support reproducibility, we provide comprehensive details regarding the inference process and evaluation protocols in Sec. 3, Sec. 4.1, and Appendix A. We are committed to releasing the complete source code, including implementation scripts and necessary dependencies, upon acceptance. This will allow other researchers to reproduce our experiments and extend ColorCtrl in future work.
@@ A.1 Inference Settings
The target object for mask extraction is determined using the “blended_word” keywords provided by PIE-Bench (Ju et al., 2024) and ColorCtrl-Bench. We use a mask threshold of $\epsilon=0.1$ which is the same as DiTCtrl, which consistently yields strong performance across SD3, FLUX.1-dev, and CogVideoX-2B. Despite being relatively coarse, this threshold is sufficient due to the strong global adaptation ability of modern generative models, which enables them to propagate edits from sparse color cues to semantically aligned regions.
For Step1X-Edit (Liu et al., 2025b) and FLUX.1 Kontext dev (Black Forest Labs, 2025), we use their official code and setting with resolution 1024 $\times$ 1024.
@@
A.2 Sampling Details
A.3 Implementation of Compared Methods
  • FireFlow (Deng et al., 2025b): In SD3, we observe that it is difficult to select a suitable end timestep for FireFlow, as setting it too high often leads to artifacts and generation failures, as shown in Fig. 9. To mitigate quality degradation caused by excessive editing steps, we restrict editing to timesteps 0 through 3 across all blocks in SD3. In contrast, FLUX.1-dev does not exhibit this issue, and using the same range (timesteps 0 to 3) yields results comparable to those reported in the original paper. For video generation, the editing ends at timestep $9$. During editing, value tokens ($\bm{V}$) are copied from the source to the target.
Figure 9: Examples of FireFlow and RF-Solver with difference end timesteps on SD3. The selected end timestep refers to the setting used in the benchmark evaluation, while the higher end timestep denotes a larger value chosen for comparison purposes.
A.4 ColorCtrl-Bench Construction
Here, we provide details on the construction of ColorCtrl-Bench. Following a similar approach to that in Ju et al. (2024), we use GPT to generate a dataset of tuples, each consisting of a source prompt, a target prompt, a subject token, and an instruction, consistent with the format used in PIE-Bench. These tuples represent an image before and after a color editing operation, along with an instruction describing the transformation and a subject token indicating the object to be modified. The exact prompting template is shown in Fig. 23.
A.5 Evaluation Details
Appendix B More Results and Analysis
B.1 Downstream Application: Synthetic Color Editing and Fine-Tuning
We generate 40k pairs of color-editing examples by applying ColorCtrl to FLUX.1-dev, and use this synthetic corpus to fine-tune Step1X-Edit-v1.1 (Liu et al., 2025b) for 2,300 H800 GPU-hours. This downstream model shows clear gains on color-editing tasks, especially under complex light–matter interactions such as reflections and refractions, while better preserving object details. Leveraging the multilingual capability of Step1X-Edit-v1.1, we also compare the original and fine-tuned models with Chinese edit instructions, as shown in Fig. 10. This highlights our potential of building pairwise color editing data in the future.
Figure 10: Step1X-Edit-v1.1 fine-tuned on ColorCtrl-generated color-editing data vs. original. Results are produced with Chinese edit instructions and English translations are shown below.
B.2 User Study
A total of 24 expert participants conducted pairwise evaluations on 34 image cases from PIE-Bench and ColorCtrl-Bench. For each comparison, presentation order placement was randomized and method identities were blinded. Raters chose the preferred result according to a predefined rubric (edit success, naturalness, background preservation). Overall, ColorCtrl was preferred in 64.71% of comparisons, outperforming all baselines (Tab. 8). The user-study interface is shown in Fig. 22.
B.3 More Results of Image Editing
Table 5: Quantitative image results compared with training-free methods on ColorCtrl-Bench. Results for FireFlow on SD3 are omitted due to consistency worse than using fixed seeds.
@@ B.3 More Results of Image Editing
Table 6: Quantitative image results compared with commercial models on ColorCtrl-Bench.
@@
Table 7: User study results.
Method Preference (%)
SDEdit (Meng et al., 2022) 0
RF-Solver (Wang et al., 2025) 0
FireFlow (Deng et al., 2025b) 0
UniEdit-Flow (Jiao et al., 2025) 0
FlowEdit (Kulikov et al., 2024) 0
DiTCtrl (Cai et al., 2025) 0
FLUX.1 Kontext Max (Black Forest Labs, 2025) 23.53 $\pm$ 14.00
GPT-4o Image Generation (OpenAI, 2025) 11.76 $\pm$ 8.31
Ours 64.71 $\pm$ 9.90
Table 8: Quantitative video results compared with baselines on ColorCtrl-Bench.
Method Canny BG Preservation $\text{Clip Similarity}$$\uparrow$
$\text{SSIM}$$\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ Whole Edited
Fix seed 0.6671 19.49 0.8489 28.01 26.66
FireFlow (Deng et al., 2025b) 0.7957 28.78 0.9323 25.26 24.12
RF-Solver (Wang et al., 2025) 0.8102 32.64 0.9562 25.36 24.16
SDEdit (Meng et al., 2022) 0.8281 34.30 0.9626 24.74 23.46
DiTCtrl (Cai et al., 2025) 0.7270 26.14 0.9165 26.54 25.14
UniEdit-Flow (Jiao et al., 2025) 0.8002 35.33 0.9613 25.39 24.14
Ours 0.8885 38.68 0.9893 27.32 26.16
Figure 11: Examples of image editing results.
B.4 More Results of Video Editing
@@
Tab. 8 shows quantitative results of video editing tasks on ColorCtrl-Bench. Fig. 19 and Fig. 20 present video editing results on ColorCtrl-Bench, while Fig. 21 shows additional results on PIE-Bench. Our method consistently outperforms all baselines, achieving superior color editing accuracy and better structural consistency across frames.
B.5 Additional Evaluation Protocol
B.5.1 Structural-Preservation Metrics
Beyond the Canny-SSIM metric used in the main paper, we report two additional gradient-based metrics to further verify edge and fine-detail consistency under strong color edits:
  • Y-Gradient Cosine Similarity (Y-GCS), where we convert both images to YUV, keep only the luminance channel $Y$, compute Sobel gradient magnitudes, normalize them by the shared global maximum, and take the cosine similarity between the two gradient maps, capturing global alignment of edge and fine-detail patterns while being largely insensitive to color changes.
  • Y-Gradient Magnitude Similarity (Y-GMS), where on the same luminance gradients we compute a per-pixel Gradient Magnitude Similarity (GMS) (Xue et al., 2013) map and average it over the image, providing a more localized and standard gradient-based IQA measure (Wang and Bovik, 2006) of structural consistency.
B.5.2 Editing-Quality Metrics
Beyond the standard CLIP similarity used in PIE-Bench, we observe that CLIP-based scores often exhibit over-saturated preferences and suffer from inaccurate attribute binding (Kang et al., 2025), making them suboptimal for disentangling the success of color-specific edits. Instead, following recent instruction-based image editing works (Liu et al., 2025b; Wu et al., 2025a; Deng et al., 2025a; Wu et al., 2025b; Cui et al., 2025), we adopt the VIEScore metrics (Ku et al., 2024) from GEdit-Bench (Liu et al., 2025b), which are originally designed to evaluate a wide range of instruction-based editing tasks, including color editing. Therefore, we directly reuse their definitions without any modification. Concretely, we use:
  • Semantic Consistency (SC), which measures whether the intended edit has been correctly applied.
  • Perceptual Quality (PQ), which reflects the naturalness of the result and the absence of artifacts or over-saturation.
Each score ranges from 0 to 10 (higher is better) and is produced by a state-of-the-art multimodal LLM evaluator, GPT-4o [^2] API access as of Nov. 2025 (Hurst et al., 2024). For video inputs, we uniformly sample three frames per clip, compute SC and PQ for each frame, and report the mean over the three frame-wise scores as the final scores for that video.
In combination, the structural-preservation metrics (Y-GCS, Y-GMS) diagnose whether edits preserve structure and details, while SC specifically evaluates whether the color edit is successful and PQ captures potential over-saturation or unnatural editing artifacts.
Table 9: Additional quantitative image results compared with training-free methods on PIE-Bench.
Method SD3 FLUX.1-dev
[…]
Fix seed 0.5430 0.6978 5.300 7.475 0.4905 0.7579 4.775 8.325
FireFlow (Deng et al., 2025b) 0.6105 0.7271 5.600 7.650 0.8020 0.9107 1.875 8.550
RF-Solver (Wang et al., 2025) 0.7339 0.7891 5.525 8.075 0.8068 0.9127 1.825 8.675
SDEdit (Meng et al., 2022) 0.7339 0.8109 1.950 7.975 0.8797 0.8967 0.350 7.900
DiTCtrl (Cai et al., 2025) 0.9074 0.9281 2.000 7.925 0.8553 0.9049 2.275 8.425
FlowEdit (Kulikov et al., 2024) 0.9027 0.9183 2.800 7.925 0.9061 0.9337 1.625 8.175
UniEdit-Flow (Jiao et al., 2025) 0.9034 0.9264 2.075 8.150 0.8320 0.9261 1.825 8.675
Ours 0.9244 0.9316 7.725 8.350 0.9555 0.9467 7.100 9.025
Table 10: Additional quantitative image results on PIE-Bench compared with commercial models on PIE-Bench.
Method SD3 FLUX.1-dev
[…]
Fix seed 0.5430 0.6978 5.300 7.475 0.4905 0.7579 4.775 8.325
FLUX.1 Kontext Max (Black Forest Labs, 2025) 0.7533 0.8104 8.750 7.393 0.5674 0.8079 8.525 7.625
GPT-4o Image Generation (OpenAI, 2025) 0.5901 0.7435 8.923 6.769 0.5547 0.7906 8.900 7.050
Ours 0.9244 0.9316 7.725 8.350 0.9555 0.9467 7.100 9.025
Table 11: Additional quantitative video results compared with baselines on PIE-Bench.
[…]
Fix seed 0.5537 0.6532 4.375 6.708
FireFlow (Deng et al., 2025b) 0.8278 0.8419 2.508 6.967
RF-Solver (Wang et al., 2025) 0.8517 0.8545 2.542 6.850
SDEdit (Meng et al., 2022) 0.9167 0.9102 0.833 6.167
DiTCtrl (Cai et al., 2025) 0.7027 0.8396 4.933 6.258
UniEdit-Flow (Jiao et al., 2025) 0.8446 0.8599 3.317 6.867
Ours 0.9383 0.9299 7.550 6.975
B.5.3 Results
Tabs. 9 and 11 report additional quantitative metrics on PIE-Bench for SD3, FLUX.1-dev, and CogVideoX-2B, comparing our method against training-free baselines. Across all settings, our approach consistently achieves the highest Y-GCS and Y-GMS scores, further highlighting its superior ability to preserve edge structure and local details under strong color edits. In contrast, the low SC scores of the other baselines indicate that they rarely perform the intended color changes correctly, which is consistent with the qualitative comparisons in the main paper. This also shows that CLIP similarity alone is insufficient to reveal our advantage specifically on color editing.
Tab. 10 compares our method with commercial models. Although they can obtain relatively high SC, their lower PQ scores support our observation that they tend to over-edit and introduce over-saturated, unrealistic edits. Their low Y-GCS and Y-GMS further demonstrate that they fail to maintain structural and textural consistency.
Overall, these results show that existing training-free methods do not handle color editing well, highlighting that accurate color editing is a challenging problem. Our method is able to achieve high-quality color changes while strongly preserving structure without training.
B.6 More Results on Real Image Editing
Table 12: Quantitative real image results compared with training-free models on PIE-Bench.
[…]
$\text{SSIM}$$\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ Whole Edited
PnpInversion (Ju et al., 2024) 0.6178 21.60 0.7785 26.10 21.41 0.8142 0.8106 3.650 6.575
FlowEdit (Kulikov et al., 2024) 0.6524 21.46 0.8324 25.71 21.05 0.7463 0.8222 3.275 7.700
Ours 0.7683 23.87 0.8890 25.88 20.52 0.9366 0.8969 5.450 8.425
Figure 12: Qualitative real image results compared with training-free models on PIE-Bench.
We further compare our method with the latest MM-DiT-based FlowEdit (Kulikov et al., 2024) and the classic U-Net–based PnP-Inversion (Ju et al., 2024) on real images from PIE-Bench. For both baselines, we observe that obtaining reasonable results requires carefully tuning which attention layers to edit and which timesteps to intervene at. The optimal choices vary across backbones (SD1.5, SDXL, SD3, FLUX.1-dev) and settings (noise-to-image vs. real-image editing), making their robustness and usability rather limited. For fairness, we therefore report results using the official implementations with their recommended default hyperparameters.
In contrast, our method can be applied to real images with exactly the same configuration as in the noise-to-image setting: we simply edit all layers and all timesteps, without any changes. To the best of our knowledge, this is the first attention-control editing method that does not require any hand-tuned hyperparameters when changing backbones or settings.
Quantitatively, Tab. 12 shows that our approach consistently and substantially outperforms both FlowEdit and PnP-Inversion on structural-preservation metrics, including Canny-SSIM, Y-GCS, and Y-GMS, as well as background-preservation metrics. While PnP-Inversion sometimes achieves higher CLIP similarity, this must be interpreted with caution in light of the discussion in Sec. B.5.2: CLIP-based scores often exhibit over-saturated preferences and cannot reliably capture artifacts introduced by edits. Once we consider SC and PQ, it becomes clear that PnP-Inversion in fact performs poorly at changing colors as instructed and often produces visually unnatural results, as illustrated in Fig. 12. Overall, our method remains state-of-the-art for real-image color editing, matching the strong structural preservation seen in the noise-to-image setting while achieving significantly better edit quality without any modification of the inference settings.
B.7 More Results of Multi-Object Editing
Figure 13: Examples of multi-object editing cases.
In addition to the first case in Fig. 1 and the last case in Fig. 11, which already demonstrate our multi-object editing ability, we provide further examples in Fig. 13. As shown, our method can robustly handle scenes with many objects, such as the first example where two separate piles of pills, each containing many pills of a different color, are edited simultaneously. In the second example, multiple objects are spatially close or even overlapping, yet our method still correctly assigns different target colors to each object. In the final example, we not only edit multiple chess pieces but also accurately update their corresponding reflections on the board, further illustrating precise multi-object editing.
B.8 Ablation Study of Mask Threshold
The threshold $\epsilon$ for mask extraction is the only hyperparameter in ColorCtrl. We conduct an ablation study in Fig. 14 to evaluate the sensitivity of our method to $\epsilon$. As shown, the performance is not very sensitive to this parameter: around our default choice $\epsilon=0.1$, using a stricter threshold ($\epsilon=0.2$) or a more relaxed one ($\epsilon=0.05$) leads to only minor visual differences. Even when we completely remove the mask extraction and color preservation modules ($\epsilon=0$), the structural details of the background remain unchanged. The main effect of using too small a threshold is a slight hue shift in some background regions, but the overall results are still visually coherent.
Figure 14: Ablation study of the mask extraction threshold $\epsilon$. For better visibility, we visualize $1-\text{mask}$.
Figure 15: Examples of additional applications enabled by reducing the preservation strength.
B.9 More Application and Ability
ColorCtrl is designed to edit color while keeping geometry material and light-matter interaction strictly fixed. However, this does not mean it can only produce extremely conservative edits. In fact, by appropriately reducing the number of effective timesteps of structure preservation (i.e., weakening the preservation strength), we can obtain a variety of different applications and effects while still preserving structure, as illustrated in Fig. 15, including style transfer, material change, texture change, and transparency change. In the first example, a real-image style is turned into a cartoon style while preserving geometry extremely well, without requiring a separately trained ControlNet (Zhao et al., 2023). In the second example, a smooth rubber pig is transformed into a metallic bronze pig, showing characteristic metallic texture and reflections while its shape remains unchanged. In the third example, rust-like textures are added to an originally brand-new bicycle. In the last example, we change the transparency of the liquid in a glass, turning clear water into opaque milk, while keeping both the glass structure and the non-edited regions well preserved.
B.10 Limitations